# Online Navigation Planning for Long-term Autonomous Operation of Underwater Gliders

### v1

@@ IV-A Problem Statement
As mentioned, glider robots execute dives and surface every few hours. They float while transmitting scientific information and receiving control instructions, after which they dive. During dives, they lack the ability to communicate or localise precisely, rendering the states between dives practically unobservable. Therefore, our model of the glider navigation planning problem considers dives as discrete events. This enables us to approach the high-level navigation planning problem over a relatively long time horizon, while the control in between dives is handled by a lower-level controller on board the glider.
@@
Let us first introduce some fundamental concepts and notation. The glider’s position is represented as $\mathbf{p}=(\lambda,\phi)$, where $\lambda$ and $\phi$ denote the longitude and latitude respectively, and $\tau$ is a continuous timestamp in seconds. A glider state is denoted as $\mathbf{s}=(\mathbf{p},\tau)$. We refer to states at MDP time $t$ as $\mathbf{s}_{t}$, and also use $\mathbf{p}_{t}$ and $\tau_{t}$ as shorthands. We are given a start state $\mathbf{s}_{0}$ and goal position $\mathbf{p}_{\text{goal}}$ and seek to find a trajectory that minimises the total duration $\Delta=\tau_{T}-\tau_{0}$ for surfacing within a radius $\rho$ of the goal at the terminal timestep $T$.
@@
Glider navigation is strongly impacted by currents. We represent them as the vector $\mathbf{q}(\lambda,\phi,z,\tau)=[u,v,w]^{\top}$ denoting the velocity in the eastward, northward, and downward components respectively. Here, $z$ represents the depth. The bathymetry $\mathbf{b}(\lambda,\phi)=z_{\text{bathy}}$ identifies the maximum ocean depth at each position, and is relevant since gliders typically inflect when approaching the ocean floor. In contrast to currents, which we cannot measure or predict fully accurately, we may assume bathymetry is known and fixed on the timescales of the considered planning problem.
@@
Each instance of $I$ of the glider navigation planning problem is therefore defined as the tuple $I=(\mathbf{p}_{0},\tau_{0},\mathbf{p}_{\text{goal}},\rho,\mathbf{q}(\cdot),\mathbf{b}(\cdot))$.
IV-B MDP Formulation
@@
To formulate the SSP MDP, we must define each element of $(\mathcal{S,A,T,C,G})$ as introduced in Section III-B. We consider each dive as a discrete decision, and therefore the MDP time $t$ advances by $1$ with each dive.
@@
Actions. Actions $\mathcal{A}$ are defined as tuples $\mathbf{a}=(\alpha,\mathbf{c})$, where $\alpha$ denotes the dive direction, while $\mathbf{c}$ is itself a tuple consisting of the dive control parameters.
We opt to discretise the dive direction component of the action space. This allows us to inject the knowledge that small increments of heading change are largely irrelevant given the noise in the system, and avoids the challenges of working directly with a continuous action space. Furthermore, we leverage the observation that glider headings $\psi$ are set relative to the bearing $\beta$ to the goal position $\mathbf{p}_{\text{goal}}$. We note the distinction between the bearing to the goal (a geometric quantity) and the vehicle heading (a control variable). Setting $\psi=\beta$ corresponds to aiming the glider straight-to-goal which, as previously mentioned, is the default mechanism by which Slocum gliders set their heading towards a distant waypoint. Instead, we may wish to set a heading that differs from the bearing in order to account for the effect of the currents. Therefore, we allow as actions a discrete set of relative bearings $\alpha=\beta-\psi$. Assuming a north-east-down reference frame, positive values of $\alpha$ correspond to diving clockwise relative to the goal, while negative values correspond to diving anti-clockwise. This preserves a lower branching factor compared to deciding the heading directly, while prioritising actions that are more likely to lead to progress towards the goal. We illustrate this in Figure 4.
@@
Transitions. Transitions $\mathcal{T}$ govern the probability of the glider surfacing in state $\mathbf{s}_{t+1}$ after diving from state $\mathbf{s}_{t}$ using action $\mathbf{a}_{t}$. In this problem, transitions are impacted by currents $\mathbf{q}$ and noise in execution of the intra-dive control policy, and cannot be modelled fully accurately. Transitions are therefore stochastic and highly uncertain. To devise navigation plans, we thus require access to a sample-based simulator that can generate samples of likely surfacing states. This component is described in depth in Section IV-C.
@@
Costs. The cost function $\mathcal{C}$ we consider is $\mathcal{C}(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1})=\tau_{t+1}-\tau_{t}$, i.e., we incur a cost equal to the duration of the dive. Alternative cost definitions may also account for the energy usage of each set of controls; however, in our considered experimental setup, this is approximately constant.
@@
Goals. The goal states $\mathcal{G}$ are defined as $\mathcal{G}=\{\mathbf{s}\ |\ d(\mathbf{p},\mathbf{p}_{\text{goal}})\leq\rho\}$, where $\rho$ is a tolerance radius and $d$ is geodesic distance.
@@
Regarding the impact of currents on transitions $\mathcal{T}$, as we cannot accurately measure and predict currents across the entire water column, we must rely on forecasts $\mathbf{\hat{q}}$. Many ocean models exist, which vary in accuracy as well as spatial and temporal resolutions. Such models employ discretisation by splitting coordinates, depth, and time into intervals. Concretely, as we focus our deployment and evaluation in the North Sea region, we make use of the AMM15 [27] model, which provides forecasts at hourly intervals, over 33 depth levels, and over grid cells of approximately 1.5km $\times$ 1.5km. We obtain these forecasts via the Met Office marine data service [28]. These forecasts have a horizon of 5 days. This model does not include vertical speeds; therefore, we assume $w=0$, i.e., $\hat{\mathbf{q}}(\lambda,\phi,z,\tau)=[u,v,0]^{\top}$.
IV-C Custom Glider Dive Simulator
@@
Our approach requires a simulator that can generate samples from the transition function $\mathcal{T}$. The simulator must balance the accuracy of sampled transitions with execution speed. As mentioned in Section III-C, the number of trials that can be ran in a fixed computational budget strongly influences MCTS performance. Constructing a search tree over a horizon of several dives is crucial for going beyond reactive current correction.
@@
The best available open source simulator for Slocum gliders is glidersim [29]. This state-of-the-art approach is highly detailed and takes into account glider actuator behaviour as well as the hydrodynamics of Slocum gliders [30, 31]. However, this comes at the expense of costly runtime. The method also does not account for the unknowns surrounding the current forecasts and it is unsuitable for use in our approach. Thus, we set out to devise a custom simulator.
@@
To design a faster simulator more appropriate for sample-based planning methods, we leverage the following observations. Firstly, the majority of large-scale, publicly available ocean models (including AMM15) do not include vertical speeds. As such, we cannot use forecasts to account for vertical influences on the glider’s depth rate, and we can treat it as approximately constant. Secondly, the glider moves in a largely steady-state (i.e., constant speed and attitude relative to the water) when performing the sawtooth motion. This is not the case at the bottom and top of each yo, when the onboard controller changes the glider pose and velocity. Nevertheless, these transient phases can be considered negligible in terms of the impact on the overall trajectory. This motivates our simulation approach that chains together steady-state segments.
@@
1:Parameters: $\Theta=\{\theta_{\text{drag\_i}},\theta_{\text{drag\_j}},$   $\theta_{\text{curr\_mag}},\theta_{\text{curr\_dir}},\theta_{\text{curr\_min}},$   $\theta_{\text{motion\_mag}},\theta_{\text{motion\_dir}},\theta_{\text{motion\_min}}\}$
@@
3:  $(\mathbf{p}^{\prime},\tau^{\prime})\leftarrow\mathbf{s}$
4:  halfYosDone $\leftarrow 0$
5:  while halfYosDone $2*\mathbf{a}.\mathbf{c}.n_{\text{yos}}$ do
6:   if halfYosDone $\text{mod}\ 2=0$ then
7:    $(\mathbf{p},\tau)\leftarrow$ SteadyDown($\mathbf{s},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b}$)
8:   else
9:    $(\mathbf{p},\tau)\leftarrow$ SteadyUp($\mathbf{s},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b}$)
10:   end if
11:   halfYosDone += 11
12:  end while
13:  $\mathbf{s}^{\prime}=(\mathbf{p},\tau)$
14:  return $\mathbf{s}^{\prime}$
15:Function SteadyDown($\mathbf{s},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b}$):
16:  $\psi\leftarrow$ HeadingFromAction($\mathbf{a}.\alpha,\beta$)
17:  $s,w_{g}\leftarrow$ GliderFlightModel($\mathbf{a}.\mathbf{c}$)
18:  $u_{g},v_{g}\leftarrow s\cdot\sin{\psi},s\cdot\cos{\psi}$
19:  $u_{g},v_{g}\leftarrow$ ApplyMotionNoise($u_{g},v_{g}$)
20:  $\mathbf{g}\leftarrow[\mathbf{p},z,\tau]^{\top}$
21:  $\dot{\mathbf{g}}\leftarrow[u_{g},v_{g},-w_{g}]^{\top}$
22:  while $\mathbf{g}[2]>\min{(\mathbf{a}.\mathbf{c}.z_{\text{bottom}},\mathbf{b}(\mathbf{g}))}$ do
@@
24:  end while
25:  $(\mathbf{p}^{\prime},\tau^{\prime})\leftarrow(\mathbf{g}[1],\mathbf{g}[3])$
26:  return $(\mathbf{p}^{\prime},\tau^{\prime})$
27:Function ToNextCell($\hat{\mathbf{q}},\mathbf{g},\dot{\mathbf{g}},z_{\text{lim}}$)
28:  $u,v,\_\leftarrow\hat{\mathbf{q}}(\mathbf{g}$)
29:  $\tilde{\mathbf{q}}(\mathbf{g})\leftarrow$ GetNoisyCurrents($u,v$)
@@
32:  $\mathbf{g}\leftarrow\mathbf{g}+\dot{\mathbf{g}}_{\text{true}}\cdot(\Delta\tau+\varepsilon)$
33:  return $\mathbf{g}$
34:Function SteadyUp($\mathbf{s},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b}$):
35:  /* Similar to SteadyDown but with $w_{g}$ in $\dot{\mathbf{g}}$ and condition $\mathbf{g}[2]<\mathbf{a}.\mathbf{c}.z_{\text{top}}$ */
36:Function TimeToNextCell($\hat{\mathbf{q}},\mathbf{g},\dot{\mathbf{g}}_{\text{true}},z_{\text{lim}}$)
37:  /* Calculates time until next cell of forecast $\hat{\mathbf{q}}$ is reached using geographical distances and glider speed $\dot{\mathbf{g}}_{\text{true}}$. */
38:Function GetNoisyCurrents($u,v$):
39:  $|q|\leftarrow\sqrt{u^{2}+v^{2}},\alpha_{\text{curr}}\leftarrow\text{atan}(v,u)$
@@
41:  $|q|$ += magN, $\alpha_{\text{curr}}$ += currN
42:  $|q|=\max(|q|,\theta_{\text{curr\_min}})$
43:  return polarToVec($|q|,\alpha_{\text{curr}}$)
44:Function ApplyMotionNoise($u,v$):
@@
45:  /* Similar to GetNoisyCurrents but with magN and dirN generated using random walks with parameters $\theta_{\text{motion\_mag}},\theta_{\text{motion\_dir}}$ and minimum speed $\theta_{\text{motion\_min}}$. */
@@
The pseudocode for our custom glider dive simulator is given in Algorithm 1. The inputs to the simulator are the glider state $\mathbf{s}$, action $\mathbf{a}$, bearing $\beta$, current forecasts $\hat{\mathbf{q}}$ and bathymetry $\mathbf{b}$. The output of the simulator is the post-dive state $\mathbf{s}^{\prime}$. We note that, unlike in the MDP states described in Section IV-B, in which the glider is always at the surface, the internal glider states in the simulator also contain a depth level. We denote this internal state as $\mathbf{g}$. Querying the forecast and bathymetry is denoted $\hat{\mathbf{q}}(\mathbf{g})$ and $\mathbf{b}(\mathbf{g})$.
@@
The main loop of the algorithm (SimulateDive) simulates the characteristic seesaw-like motion by chaining together upwards and downwards steady-state phases (SteadyUp and SteadyDown) until the number of yos to perform is exhausted. The SteadyDown method simulates a downward steady-state phase by repeatedly propagating the glider to the next forecast cell. The SteadyUp method does the same except with the glider moving upwards. The general idea is that, as the forecast currents are constant in a grid cell, the glider state can be propagated linearly within that cell until the cell border is hit (TimeToNextCell). Glider horizontal speed $s\in\mathbb{R}^{+}$ and vertical speed (depth rate) $w_{g}\in\mathbb{R}^{+}$ are calculated (GliderFlightModel) using the approach of Merckelbach et al. [31]. This takes into account the control parameters in $\mathbf{c}$ as well as water properties. The flight model parameters are fit to historical glider data from the region of interest. Acceleration is applied to the glider speeds as determined by the current forecast and drag coefficients $\theta_{\text{drag\_i}},\theta_{\text{drag\_j}}$ that are part of the simulator parameters.
@@
By default, transitions $\mathcal{T}$ generated by this simulator would be deterministic. However, as we would like the produced navigation plans to be robust to forecast and control errors, we inject two types of noise in order to yield several plausible samples of stochastic transitions given the same inputs. Namely, we inject current noise and motion noise.
@@
We use a simple Gaussian noise model for the current noise (GetNoisyCurrents). Independent noise is added to the magnitude and direction by sampling from Gaussian distributions with mean 0 and configurable standard deviations $\theta_{\text{curr\_mag}}$, $\theta_{\text{curr\_dir}}$. These perturbations are applied consistently for all forecasts $\hat{\mathbf{q}}$ given a particular random seed, corresponding to a systematic bias in the forecast.
@@
To implement motion noise (ApplyMotionNoise), we update the glider velocity with a random walk to reflect the noise of the internal dead-reckoning characterised by a very inaccurate state estimation and other local perturbations – a noise model suggested by glider pilots. We simulate a symmetric random walk over integers with absorbing barriers for which the probability of the value decreasing, staying put, or increasing is uniform. Separate walks are used for the magnitude and direction. The values of the random walk states are multiplied with the $\theta_{\text{motion\_mag}}$ and $\theta_{\text{motion\_dir}}$ noise parameters then added to the glider magnitude and direction.
@@
The simulator parameters $\Theta$, enumerated in Line 1 of Algorithm 1, control the impact of the currents on glider speed and the amount of noise that is injected into the simulator. As these parameters are challenging to establish a priori and depend significantly on the deployment vehicle, location, and conditions, we propose a procedure to tune them using past glider data. We describe this procedure in the next section.
IV-D Simulator Parameter Tuning

### v2

@@ 4.A Problem Statement
As mentioned, gliders execute dives and surface every few hours. They float while transmitting scientific information and receiving control instructions, after which they dive. During dives, they lack the ability to communicate or localise precisely, rendering the states between dives practically unobservable. Therefore, our model of the problem considers surfacings as discrete events. This enables us to approach the high-level navigation planning problem over a relatively long time horizon, while the control in between dives is handled by a lower-level controller on board the glider.
@@
The glider’s position is represented as $\mathbf{p}=(\lambda,\phi)$, where $\lambda$ and $\phi$ denote the longitude and latitude respectively. A glider state is denoted as $\mathbf{s}=(\mathbf{p},\tau)$, where $\tau$ is a continuous timestamp in seconds. We refer to states at MDP time $t$ as $\mathbf{s}_{t}$, and also use $\mathbf{p}_{t}$ and $\tau_{t}$ as shorthands. We are given a start state $\mathbf{s}_{0}$ and goal position $\mathbf{p}_{\text{goal}}$ and seek to find a trajectory that minimises the total duration ${\Delta\tau}_{\text{total}}=\tau_{T}-\tau_{0}$ for surfacing within a radius $\rho$ of the goal at the terminal timestep $T$.
@@
Glider navigation is strongly impacted by currents. We represent them as the vector $\mathbf{q}(\lambda,\phi,z,\tau)=[u,v,w]^{\top}$ denoting the velocity in the eastward, northward, and downward components respectively. Here, $z$ represents the depth. The bathymetry $\mathbf{b}(\lambda,\phi)=z_{\text{bathy}}$ identifies the maximum ocean depth at each position and is relevant since gliders typically inflect when approaching the ocean floor. In contrast to currents, which we cannot measure or predict fully accurately, we may assume bathymetry is known and fixed on the timescales of the considered planning problem.
@@
Each instance $I$ of the glider navigation planning problem is therefore defined as the tuple $I=(\mathbf{p}_{0},\tau_{0},\mathbf{p}_{\text{goal}},\rho,\mathbf{q}(\cdot),\mathbf{b}(\cdot))$.
4.B Markov Decision Process Formulation
MDPs are a widely employed formalisation of decision-making [22, 23]. We define an MDP as the tuple $(\mathcal{S,A,T,C,G})$, where $\mathcal{S}$ is the state space defining all configurations the agent may find itself in, $\mathcal{A}$ is the action space defining valid controls, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ denotes the transition function satisfying $\sum_{a\in\mathcal{A}}{\sum_{s^{\prime}\in\mathcal{S}}{\mathcal{T}(s,a,s^{\prime})}}=1\ \forall s\in\mathcal{S}$, and $\mathcal{C}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,\infty)$ is a cost function. Also let $\mathcal{G}\subseteq\mathcal{S}$ denote a set of absorbing goal states s.t. $\forall s_{g}\in\mathcal{G}\land s^{\prime}\notin\mathcal{G}$ we have $\mathcal{T}(s_{g},a,s_{g})=1$, $\mathcal{T}(s_{g},a,s^{\prime})=0$, and $\mathcal{C}(s_{g},a,s_{g})=0$. This class of models is known as stochastic shortest-path (SSP) MDPs. A deterministic, stationary, Markovian policy is a mapping $\pi:\mathcal{S}\to\mathcal{A}$. The value function of such a policy for a state $s$ is defined as $V^{\pi}(s)=\mathbb{E}[\sum_{t^{\prime}=0}^{\infty}{C_{t^{\prime}+t}^{\pi_{s}}}]$, where $C_{t^{\prime}+t}^{\pi_{s}}$ is a random variable denoting the cost incurred as a result of executing policy $\pi$ starting in state $s$ for all $t$. We seek to find the optimal policy $\pi^{*}$ which satisfies $V^{*}(s)=\min_{\pi}V^{\pi}(s)$, i.e., the policy that achieves the minimal cost in expectation at each state.
@@ 4.B Markov Decision Process Formulation
To formulate glider navigation planning as an SSP MDP, we must define each element of $(\mathcal{S,A,T,C,G})$. We consider each dive as a discrete decision, hence the MDP time $t$ advances by $1$ with each dive.
@@
Actions. Actions $\mathcal{A}$ are defined as tuples $\mathbf{a}=(\alpha,\mathbf{c})$, where $\alpha$ denotes the dive direction, while $\mathbf{c}$ is itself a tuple consisting of the dive control parameters. We opt to discretise the dive direction component of the action space. This allows us to incorporate the knowledge that small increments of change in heading are largely irrelevant given the noise in the system, and avoids the challenges of working directly with a continuous action space. Furthermore, we leverage the observation that glider headings $\psi$ are set relative to the bearing $\beta$ to the goal position $\mathbf{p}_{\text{goal}}$, as illustrated in Figure 6. We note the distinction between the bearing to the goal (a geometric quantity) and the vehicle heading (a control variable). Setting $\psi=\beta$ corresponds to aiming the glider straight-to-goal which, as previously mentioned, is the default mechanism by which Slocum gliders set their heading towards a distant waypoint. Instead, we may wish to set a heading that differs from the bearing in order to account for the effect of the currents. Therefore, we allow as actions a discrete set of relative bearings $\alpha=\beta-\psi$. Assuming a north-east-down reference frame, positive values of $\alpha$ correspond to diving clockwise relative to the goal, while negative values correspond to diving anti-clockwise. This preserves a lower branching factor compared to deciding the heading directly, while prioritising actions that are more likely to lead to progress towards the goal.
@@
Transitions. Transitions $\mathcal{T}$ govern the probability of the glider surfacing in state $\mathbf{s}_{t+1}$ after diving from state $\mathbf{s}_{t}$ using action $\mathbf{a}_{t}$. In this problem, transitions are impacted by uncertain currents $\mathbf{q}$ and glider motion, and cannot be modelled fully accurately. Transitions are therefore stochastic and highly uncertain. To devise navigation plans, we thus require access to a sample-based simulator that can generate samples of likely surfacing states. This component is described in depth in Section 4.C.
@@
Costs. The cost function $\mathcal{C}$ we consider is $\mathcal{C}(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1})=\tau_{t+1}-\tau_{t}$, i.e., we incur a cost equal to the duration of the dive. Alternative cost definitions may also account for the energy usage of each set of controls; however, in our considered experimental setup, this is approximately constant with respect to time.
@@
Goals. The goals $\mathcal{G}$ are defined as $\mathcal{G}=\{\mathbf{s}\in\mathcal{S}\ |\ d(\mathbf{p},\mathbf{p}_{\text{goal}})\leq\rho\}$, where $\rho$ is a tolerance radius and $d$ is geodesic distance.
@@
Regarding the impact of currents on transitions $\mathcal{T}$, as we cannot accurately measure and predict currents across the entire water column, we must rely on forecasts $\mathbf{\hat{q}}$. Many ocean models exist, which vary in accuracy as well as spatial and temporal resolutions. Such models employ discretisation by splitting coordinates, depth, and time into intervals. Concretely, as we focus our deployment and evaluation in the North Sea region, we make use of the AMM15 [24] model, which provides forecasts at hourly intervals over grid cells of $1.5\times 1.5$ km and 33 levels whose coarseness increases with depth. We obtain these forecasts via the Met Office marine data service [25]. These forecasts have a horizon of 5 days. This model does not make vertical velocities available in the publicly available outputs; therefore, we assume $w=0$, i.e., $\hat{\mathbf{q}}(\lambda,\phi,z,\tau)=[u,v,0]^{\top}$, which is a reasonable assumption in tidally dominated shelf seas.
4.C Custom Glider Dive Simulator
@@
Our approach requires a simulator that can generate samples from the transition function $\mathcal{T}$. Recall that, as described in Section 2, control instructions are decided within a fixed time budget of 5-10 minutes while the glider is floating at the surface. Given that the number of trials that can be run in a fixed time budget strongly influences MCTS performance [26], the simulator must balance the accuracy of sampled transitions with execution speed. Constructing a search tree over a horizon of several dives is crucial for going beyond reactive current correction.
@@
The state-of-the-art open source simulator for Slocum gliders is glidersim [27], whose detailed approach takes into account the actuator behaviour and hydrodynamics of Slocum gliders [28, 29]. However, this comes at the expense of costly runtime. The method also does not account for the unknowns surrounding the current forecasts and therefore it is unsuitable for use in our approach. Thus, we set out to devise a custom simulator that is more appropriate for sample-based planning methods.
@@
To do so, we leverage the following observations. Firstly, the majority of large-scale, publicly available ocean models (including AMM15) do not include vertical velocities in the published model outputs. As such, we cannot use forecasts to account for changes in the glider’s descent and ascent rate, and we can treat them as approximately constant. While vertical density gradients (stratification) can also influence vertical motion in practice, capturing these effects would require a more detailed hydrodynamic model, which we leave to future work. Secondly, the glider moves in a largely steady state (i.e., constant speed and attitude relative to the water) when performing the sawtooth motion. This is not the case at the bottom and top of each yo, when the onboard controller changes the glider pose and velocity. Nevertheless, these transient phases can be considered negligible in terms of the impact on the overall trajectory, as demonstrated by prior studies [30, 31]. This motivates our simulation approach that chains together steady-state segments.
@@
1:Params: $\Theta=\{\theta_{\text{drag\_i}},\theta_{\text{drag\_j}},\theta_{\text{curr\_mag}},\theta_{\text{curr\_dir}},\theta_{\text{curr\_min}},$        $\theta_{\text{motion\_mag}},\theta_{\text{motion\_dir}},\theta_{\text{motion\_min}}\}$
@@
3:  $\mathbf{s}^{\prime}\leftarrow\mathbf{s}$, halfYosDone $\leftarrow 0$
4:  while halfYosDone $2*\mathbf{a}.\mathbf{c}.n_{\text{yos}}$ do
5:   if halfYosDone $\text{mod}\ 2=0$ then
6:    $\mathbf{s}^{\prime}\leftarrow$ SteadyDown($\mathbf{s}^{\prime},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b}$)
7:   else
8:    $\mathbf{s}^{\prime}\leftarrow$ SteadyUp($\mathbf{s}^{\prime},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b}$)
9:   halfYosDone += 11
10:  return $\mathbf{s}^{\prime}$
11:Function SteadyDown($\mathbf{s},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b}$):
12:  $\psi\leftarrow$ HeadingFromAction($\mathbf{a}.\alpha,\beta$)
13:  $s,w_{g}\leftarrow$ GliderFlightModel($\mathbf{a}.\mathbf{c}$)
14:  $u_{g},v_{g}\leftarrow s\cdot\sin{\psi},s\cdot\cos{\psi}$
15:  $u_{g},v_{g}\leftarrow$ ApplyMotionNoise($u_{g},v_{g}$)
16:  $\mathbf{g}\leftarrow[\mathbf{p},z,\tau]^{\top}$
17:  $\dot{\mathbf{g}}\leftarrow[u_{g},v_{g},-w_{g}]^{\top}$
18:  while $\mathbf{g}[2]>\min{(\mathbf{a}.\mathbf{c}.z_{\text{bottom}},\mathbf{b}(\mathbf{g}))}$ do
@@
20:  $(\mathbf{p}^{\prime},\tau^{\prime})\leftarrow(\mathbf{g}[1],\mathbf{g}[3])$
21:  return $(\mathbf{p}^{\prime},\tau^{\prime})$
22:Function ToNextCell($\hat{\mathbf{q}},\mathbf{g},\dot{\mathbf{g}},z_{\text{lim}}$)
23:  $u,v,\_\leftarrow\hat{\mathbf{q}}(\mathbf{g}$)
24:  $\tilde{\mathbf{q}}(\mathbf{g})\leftarrow$ GetNoisyCurrents($u,v$)
@@
27:  $\mathbf{g}\leftarrow\mathbf{g}+\dot{\mathbf{g}}_{\text{true}}\cdot(\Delta\tau+\varepsilon)$
28:  return $\mathbf{g}$
29:Function SteadyUp($\mathbf{s},\mathbf{a},\beta,\hat{\mathbf{q}},\mathbf{b}$):
30:  /* Similar to SteadyDown but with $w_{g}$ in $\dot{\mathbf{g}}$ and condition $\mathbf{g}[2]<\mathbf{a}.\mathbf{c}.z_{\text{top}}$ */
31:Function TimeToNextCell($\hat{\mathbf{q}},\mathbf{g},\dot{\mathbf{g}}_{\text{true}},z_{\text{lim}}$)
32:  /* Calculates time until next forecast cell $\hat{\mathbf{q}}$ is reached using geographical distances and glider speed $\dot{\mathbf{g}}_{\text{true}}$. */
33:Function GetNoisyCurrents($u,v$):
34:  $|q|\leftarrow\sqrt{u^{2}+v^{2}},\alpha_{\text{curr}}\leftarrow\text{atan}(v,u)$
@@
36:  $|q|$ += magN, $\alpha_{\text{curr}}$ += currN
37:  $|q|=\max(|q|,\theta_{\text{curr\_min}})$
38:  return polarToVec($|q|,\alpha_{\text{curr}}$)
39:Function ApplyMotionNoise($u,v$):
@@
40:  /* Similar to GetNoisyCurrents but with generatingmagN, dirN using random walks with parameters $\theta_{\text{motion\_mag}},\theta_{\text{motion\_dir}}$ and minimum speed $\theta_{\text{motion\_min}}$. */
@@
The pseudocode for our glider dive simulator is given in Algorithm 1. The inputs to the simulator are the glider state $\mathbf{s}$, action $\mathbf{a}$, bearing $\beta$, current forecasts $\hat{\mathbf{q}}$ and bathymetry $\mathbf{b}$. The output of the simulator is the post-dive state $\mathbf{s}^{\prime}$. We note that, unlike in the MDP states described in Section 4.B, in which the glider is always at the surface, the internal glider states in the simulator also contain a depth level. We denote this internal state as $\mathbf{g}$. Querying the forecast and bathymetry are denoted $\hat{\mathbf{q}}(\mathbf{g})$ and $\mathbf{b}(\mathbf{g})$.
@@
The main loop of the algorithm (SimulateDive) simulates the characteristic seesaw-like motion by chaining together upwards and downwards steady-state phases (SteadyUp and SteadyDown) until the number of yos to perform is exhausted. The SteadyDown method simulates a downward steady-state phase by repeatedly propagating the glider to the next forecast cell. The SteadyUp method does the same except with the glider moving upwards. The general idea is that, as the forecast currents are representative for a whole grid cell, the glider state can be propagated linearly within that cell until the cell border is hit (TimeToNextCell). Glider horizontal speed $s\in\mathbb{R}^{+}$ and vertical speed (rate of change of depth) $w_{g}\in\mathbb{R}^{+}$ are calculated (GliderFlightModel) using the approach of Merckelbach et al. [29]. This takes into account the control parameters in $\mathbf{c}$ as well as water properties. The flight model parameters are fit to historical glider data from the region of interest. Acceleration is applied to the glider speeds as determined by the current forecast and drag coefficients $\theta_{\text{drag\_i}},\theta_{\text{drag\_j}}$ that are part of the simulator parameters.
@@
By default, transitions $\mathcal{T}$ generated by this simulator would be deterministic. However, as we would like the produced navigation plans to be robust to uncertainty in the forecast and glider motion, we inject current noise and motion noise in order to yield several plausible samples of transitions given the same inputs.
@@
We use a simple Gaussian noise model for the current noise (GetNoisyCurrents). Independent noise is added to the magnitude and direction by sampling from Gaussian distributions with mean 0 and configurable standard deviations $\theta_{\text{curr\_mag}}$, $\theta_{\text{curr\_dir}}$. These perturbations are applied consistently for all forecasts $\hat{\mathbf{q}}$ upon the initialisation of the simulator with a particular random seed, corresponding to a systematic bias in the forecast.
@@
To implement motion noise (ApplyMotionNoise), we update the glider velocity with a noise model suggested by glider pilots. Namely, we use a random walk to reflect the noise of the internal dead-reckoning characterised by a very inaccurate state estimation and other local perturbations. We simulate a symmetric random walk over integers with absorbing barriers for which the probability of the value decreasing, staying put, or increasing is uniform. Separate walks are used for the magnitude and direction. The values of the random walk states are multiplied with the $\theta_{\text{motion\_mag}}$ and $\theta_{\text{motion\_dir}}$ noise parameters then added to the glider magnitude and direction.
@@
The simulator parameters $\Theta$, enumerated in Line 1 of Algorithm 1, control the impact of the currents on glider speed and the amount of noise that is injected into the simulator. As these parameters are challenging to establish a priori and depend significantly on the deployment vehicle, location, and conditions, we propose a procedure to tune them using past glider data.
4.D Simulator Parameter Tuning